# Language Models Trained to do Arithmetic Predict Human Risky and Intertemporal Choice

### v1

@@ Abstract
The observed similarities in the behavior of humans and Large Language Models (LLMs) have prompted researchers to consider the potential of using LLMs as models of human cognition. However, several significant challenges must be addressed before LLMs can be legitimately regarded as cognitive models. For instance, LLMs are trained on far more data than humans typically encounter, and may have been directly trained on human data in specific cognitive tasks or aligned with human preferences. Consequently, the origins of these behavioral similarities are not well understood. In this paper, we propose a novel way to enhance the utility of LLMs as cognitive models. This approach involves (i) leveraging computationally equivalent tasks that both an LLM and a rational agent need to master for solving a cognitive problem and (ii) examining the specific task distributions required for an LLM to exhibit human-like behaviors. We apply this approach to decision-making – specifically risky and intertemporal choice – where the key computationally equivalent task is the arithmetic of expected value calculations. We show that an LLM pretrained on an ecologically valid arithmetic dataset, which we call Arithmetic-GPT, predicts human behavior better than many traditional cognitive models. Pretraining LLMs on ecologically valid arithmetic datasets is sufficient to produce a strong correspondence between these models and human decision-making. Our results also suggest that LLMs used as cognitive models should be carefully investigated via ablation studies of the pretraining data.
@@ 1 Introduction
Figure 1: (A) Pre-training and evaluation pipelines. We begin by generating a synthetic dataset comprised of mathematical equations including addition, subtraction, multiplication, and exponentiation. Arithmetic-GPT was pretrained on this synthetic dataset. After training, we froze model weights and extracted embeddings from the pretrained model, which then processes stylized choice tasks as input. These embeddings were subsequently compared with human choice data to evaluate their correspondence. (B) Ecological distributions of probabilities and values. In the top panel, English probability-describing phrases (black bars) can be modeled using a Beta(0.27,0.27)0.270.27(0.27,0.27)( 0.27 , 0.27 ) distribution. In the bottom panel, the value distribution of debits from UK bank accounts (scatterpoints) follows a power-law distribution. Figures were adapted from [47] and [38].
Scientists studying the behavior of Large Language Models (LLMs) in cognitive tasks typically performed by humans have found substantial evidence that LLMs produce performance similar to that of human participants [4, 18, 46, 9, 13, 27, 43, 8]. Like humans, LLMs often make judgments and decisions that deviate from rational norms [4, 18, 46, 8]. For instance, GPT-3 demonstrates human-like biases in risky choice, such as risk aversion and loss aversion [4], and the statistical properties of probability judgments generated by LLMs align qualitatively with those of humans [46]. LLMs also make errors in other related settings; for example, when trained on the task of predicting the next token in sequences involving arithmetic operations, they fail to learn precise arithmetic operands and instead approximate the correct results up to a certain input range [31, 24].
In this paper, we focus on risky and intertemporal choice as exemplar domains for comparing the behavior of LLMs and humans (see Figure 1). Central to both domains is the computational-level challenge of calculating expectations. To assess the benefits of engaging in a gamble, an intelligent system must be able to calculate the expected value (EV) of the gamble, typically represented as:
@@
where the value $x_{t}$ is realized at time $t$ and is discounted by a factor of $d$, reflecting the time preference of the decision-maker. Note that a risk-neutral and time-consistent agent should always select the option that maximizes EV and PV. However, extensive research in economics and psychology demonstrates that people systematically deviate from this maximizer model [21, 16, 22, 47].
Pretrained, off-the-shelf LLMs, such as the GPT and LLaMA series, have demonstrated behavioral similarities to humans in tasks involving risky and intertemporal choices [18, 4, 26]. However, the embeddings generated by these LLMs are not by default able to account for human data. For example, embeddings from the LLaMA-1-65B model, when not finetuned on human risky choices, poorly predict those choices [3]. Therefore, understanding what enables LLMs to exhibit human-like decision-making behavior remains an unresolved challenge.
We propose a hypothesis about how human-like decision patterns might be produced in LLMs for risky and intertemporal choice: such human-like behaviors might arise from an LLM trained on ecologically valid calculations of expectations. As a corollary, this hypothesis also implies that deviations from rational choice in humans could be primarily explained by computational errors during the EV or PV calculations. To test this hypothesis, we generate a series of synthetic datasets that contain expected value computations; examples include 0.5*100=+50, 0.8*1+0.8^2*10=+7.2, and 30*0.79-261*0.83=-192.93. Subsequently, we train a small, randomly-initialized language model (approximately 10M parameters) on these datasets. After training, we extract embeddings from the now pretrained language model and analyze how well they can account for human choices.
@@ 2 Background
The question of whether LLMs can serve as models of human cognition has sparked intense debate within the fields of machine learning and cognitive science [13, 30, 17, 18]. Although there are behavioral similarities between LLMs and humans, these do not inherently qualify LLMs as effective cognitive models (c.f. the Clever Hans effect) [36]. There are compelling reasons why current LLMs may not be suitable as cognitive models. First, LLMs are trained on datasets vastly larger than those available to human learners [12]. Second, LLMs may have already been trained on the test questions, particularly if the training data is undisclosed and poorly controlled [20]. Third, the inclusion of value alignment steps, such as Reinforcement Learning from Human Feedback [48] and Direct Preference Optimization [35], may artificially enhance human-like behaviors in leading LLMs. Finally, the immense size of deep neural networks and the proprietary nature of leading LLMs hinder detailed investigation into their internal representations.
@@
Researchers have taken steps to address some of the concerns associated with using LLMs as cognitive models. One branch of research looks at the potential benefits of fine-tuning off-the-shelf LLMs to better understand human cognition. For instance, fine-tuning the LLaMA-1-65B model [39] on datasets of human choices results in a model that can predict human data more accurately than traditional cognitive models [3]. Although the specific mechanisms underlying the finetuned LLM’s ability to replicate human choices remain unclear, this work serves as a proof-of-concept and suggests that the embeddings learned from extensive pretraining on Internet text and/or from the value alignment process may offer valuable insights into human cognitive processes that complement those provided by traditional cognitive models.
@@
Another line of research emphasizes the importance of the composition of synthetic datasets, which are critical for enhancing certain capabilities of LLMs [24]. These studies typically assess LLMs based on their problem-solving abilities rather than their human-likeness. For example, it has been found that larger language models tend to perform arithmetic tasks, such as addition and multiplication, better than their smaller counterparts [45]. Moreover, the order of input and the inclusion of intermediate information about task decomposition have been shown to facilitate chain-of-thought reasoning, which significantly helps smaller language models in mastering arithmetic tasks [24].
@@
Finally, there is a precedent for the idea that pretraining machine learning models on synthetic datasets can improve performance in predicting human decisions. Bourgin et al. [5] showed that a model pretrained on choice data generated from a psychological theory could perform extremely well in predicting human decisions when fine-tuned with a small amount of human data. Our approach builds on this idea, but reduces it to the most primitive components – rather than pretraining on data generated from a psychological theory, we pretrain on a task that captures the basic computations required to make rational decisions.
@@ 3 Arithmetic-GPT: A Small Language Model Trained to Perform Arithmetic
In this paper, we confront the challenges of making LLMs as cognitive models head-on. First, we define a data generation algorithm to produce synthetic datasets, thereby gaining complete control over the training data for LLMs and addressing issues related to data gaps and contamination. Second, we have direct access to the neural activation patterns that are crucial for decision-making processes. This approach allows us to more thoroughly evaluate and understand the capabilities and limitations of LLMs.
@@ 3.1 Model Details
Our small LM employs a standard architecture for a Generative Pretrained Transformer (GPT) model [41, 34]. Detailed specifications of the model architecture are provided in Table 1.
@@ 3.2 Synthetic Datasets
At the heart of the EV and PV computations lies the multiplication of two real numbers. Typically, each outcome appears in the computation as either a probability multiplied by a value, or a discount factor multiplied by a value. Probabilities are real numbers ranging from 0 to 1, represented with a precision of up to two decimal places. Similarly, values are real numbers that range from 0 to 300, also with a maximum of two decimal places. We selected these ranges to align with the numerical scope of human experimental studies. A single training example involves either the addition or subtraction of two simulated outcomes, which together constitute the left-hand side of an equation. We then compute the corresponding result and place it on the right-hand side of the equation. In total, we simulated 1M such equations.
In our experiment, we evaluate four variants of the data generation algorithm by manipulating the frequency of probabilities and values (see Table 3). The Uniform synthetic data generates probabilities and values with maximum uncertainty; probabilities are uniformly distributed between 0 and 1 (i.e., $U[0,1]$), and values range from 0 to 300 (i.e., $U[0,300]$). Conversely, the Ecological synthetic data generates probabilities following a Beta distribution $\text{Beta}(0.27,0.27)$ [47] and values according to a power-law distribution with an exponent of $-0.945$ for the same range [38]. These distributions are chosen because they have been shown to closely match the natural frequencies of probabilities and values in real-world scenarios [38, 47] (see Figure 1B for details). For both uniform and ecological synthetic datasets, we also created a matching dataset where the answers on the right-hand side are generated with a 50% chance of displaying the incorrect sign (i.e., the ablated variants in Table 3). In each of the four synthetic datasets, we randomly masked 10% of the probability values using a special <AMB> token to denote unknown probabilities.
@@ 3.3 Pretraining Details
We trained our models from scratch with a context length of 26. Batch size was set to 2048 lines of mathematical equations, randomly sampled from the synthetic datasets, and the learning rate was $10^{-3}$. To optionally terminate the training process, we designated 90% of the synthetic dataset as the training set, reserving the remaining 10% for validation. We used cross-entropy loss to measure the discrepancy between the predicted and target sequences. Training was stopped when the validation loss plateaued, with validation loss evaluated every 100 epochs. The AdamW optimizer was used.
@@ 3.4 Human Targets
To investigate whether Arithmetic-GPT contains information pertinent to explaining human decision-making, we reanalyzed recent experiments in which people were asked to make risky and intertemporal choices [33, 11, 14, 1]. The primary reason for examining these particular types of human choices is that calculations of expected value are integral to making rational decisions (see Equations 1 and 2). In other words, they are computationally equivalent tasks under the assumption of rationality.
As summarized in Table 2, we sampled four existing datasets from the literature, which included two large-scale experiments [33, 1]. In experiments involving risky choices [33, 11], participants were often presented with two options, each fully describing the details of a gamble (see Figure 1A risky choices). In cases involving ambiguous gambles where probabilities are unknown, we used the special token <AMB> to denote gambles with unknown probabilities. We excluded decision from experience trials, as these would require additional assumptions about how individuals respond to feedback. In intertemporal choice tasks [14, 7, 1], participants were also presented with two options, typically offering a choice between a smaller, sooner payoff and a larger, later payoff (see Figure 1A intertemporal choices). Without loss of generality, we fixed the annual discount factor at $d_{\text{year}}=0.85$ throughout the paper. This also corresponds to a monthly discount factor of $d_{\text{month}}=0.98$ and a daily discount factor of $d_{\text{day}}=0.99$. We rescaled the values in both options by the same factor to fit within the specified range.
@@
[33] choice13k Risky choices 15,153 13,006
[11] cpc18 Risky choices 446 270
[14] gershman20 Intertemporal choices 221 4,794
[1] agrawal23 Intertemporal choices 12,906 9,853
Note. In our analysis of the choice13k and cpc18 datasets, we excluded risky choices made with feedback (i.e., decision-from-experience trials). These trials require additional cognitive mechanisms reflecting how individuals respond to feedback, which are beyond the scope of this work.
@@ 4 Other Models
We also conduct a model comparison on human data, evaluating the following approaches: (i) classical behavioral models such as Cumulative Prospect Theory (CPT) [21] and the hyperbolic discounting model [22]; (ii) a neural network directly trained on the human datasets; (iii) an untrained Arithmetic-GPT model; and (iv) an off-the-shelf, open-weight LLM, LLaMA-3-70B-Instruct.  [^1] https://llama.meta.com/llama3/
Classical behavioral models. To explain human risky choices, CPT proposes an S-shaped utility function that is concave for gains, convex for losses, and steeper for losses than for gains (reflecting loss aversion). The utility function is represented mathematically as:
@@
Contrary to the consistent risk preferences implied by Equation 1 or other monotonic transformations of value, CPT suggests that risk preferences are inconsistent across different values and probabilities. This inconsistency results in incoherent choices when individuals are faced with risky decisions [40].
@@
where $x_{t}$ is the value to be received at future time $t$, and $k$ is the discount factor that quantifies the degree of time preference. In contrast to the consistent time preferences implied by Equation 2, the hyperbolic discounting model suggests that time preferences are inconsistent across different time horizons, leading to stronger preference to immediate over future rewards [22].
MLPs directly trained on human choices. We also implemented Multilayer Perceptrons (MLPs) with a single hidden layer containing 320 neurons and using the sigmoid activation function. These MLPs were directly trained on each of the four human datasets, using the stimuli features as input to predict choice rates. These MLPs potentially capture an upper bound of the explainable variance within human data [2]. However, they may have overlooked significant constraints from the original psychological experiments, as these models use task features as input rather than the actual stimuli presented in texts or the stylized representations required for rational agents, as in our Arithmetic-GPT.
@@
Given these features of LLaMA3, we presented each choice problem in two different formats: text-based and arithmetic equation-based. The text format converts each choice problem into a descriptive narrative, simulating the stimuli presented to human participants (see Appendix A for a detailed description of the text prompts). We instructed the model to report its selection between the two options, using the log probability of the chosen option to determine the model’s predicted choice rates for the corresponding option. In contrast, the arithmetic-equation format presents the choice problems as a series of arithmetic computations required by a rational agent. This format is identical to the input used for Arithmetic-GPT. We obtained the embeddings from LLaMA3 for each choice problem represented in the arithmetic-equation format.

### v2

@@ Abstract
The observed similarities in the behavior of humans and Large Language Models (LLMs) have prompted researchers to consider the potential of using LLMs as models of human cognition. However, several significant challenges must be addressed before LLMs can be legitimately regarded as cognitive models. For instance, LLMs are trained on far more data than humans typically encounter, and may have been directly trained on human data in specific cognitive tasks or aligned with human preferences. Consequently, the origins of these behavioral similarities are not well understood. In this paper, we propose a novel way to enhance the utility of language models as cognitive models. This approach involves (i) leveraging computationally equivalent tasks that both a language model and a rational agent need to master for solving a cognitive problem and (ii) examining the specific task distributions required for a language model to exhibit human-like behaviors. We apply this approach to decision-making – specifically risky and intertemporal choice – where the key computationally equivalent task is the arithmetic of expected value calculations. We show that a small language model pretrained on an ecologically valid arithmetic dataset, which we call Arithmetic-GPT, predicts human behavior better than many traditional cognitive models. Pretraining language models on ecologically valid arithmetic datasets is sufficient to produce a strong correspondence between these models and human decision-making. Our results also suggest that language models used as cognitive models should be carefully investigated via ablation studies of the pretraining data.
@@ 1 Introduction
Figure 1: (A) Pre-training and evaluation pipelines. We began by generating a synthetic dataset comprised of mathematical equations including addition, subtraction, multiplication, and exponentiation. Arithmetic-GPT was pretrained on this synthetic dataset. After training, we froze model weights and extracted embeddings from the pretrained model, which then processes stylized choice tasks as input. These embeddings were subsequently compared with human choice data to evaluate their correspondence. (B) Ecological distributions of probabilities and values. In the top panel, English probability-describing phrases (black bars) can be modeled using a Beta(0.27,0.27)0.270.27(0.27,0.27)( 0.27 , 0.27 ) distribution. In the bottom panel, the value distribution of debits from UK bank accounts (scatterpoints) follows a power-law distribution. Figures adapted from Zhu et al. (2020) and Stewart et al. (2006).
Scientists studying the behavior of Large Language Models (LLMs) in cognitive tasks typically performed by humans have found substantial evidence that LLMs produce performance similar to that of human participants (Binz & Schulz, 2023b; Horton, 2023; Zhu & Griffiths, 2024; Dasgupta et al., 2022; Frank, 2023b; Marjieh et al., 2023; Webb et al., 2023; Coda-Forno et al., 2024). [^1] LLMs can also display significant deviations from human behavior (e.g., Chen et al., 2023; Hagendorff et al., 2023). Like humans, LLMs often make judgments and decisions that deviate from rational norms (Binz & Schulz, 2023b; Horton, 2023; Zhu & Griffiths, 2024; Coda-Forno et al., 2024). For instance, GPT-3 demonstrates human-like biases in risky choice, such as risk aversion and loss aversion (Binz & Schulz, 2023b), and the statistical properties of probability judgments generated by LLMs align qualitatively with those of humans (Zhu & Griffiths, 2024). LLMs also make errors in other related settings; for example, when trained on the task of predicting the next token in sequences involving arithmetic operations, they fail to learn precise arithmetic operands and instead approximate the correct results up to a certain input range (Nogueira et al., 2021; Lee et al., 2023).
In this paper, we focus on risky and intertemporal choice as exemplar domains for comparing the behavior of language models and humans (see Figure 1). Central to both domains is the computational challenge of calculating expectations. To assess the benefits of engaging in a gamble, an intelligent system must be able to calculate the expected value (EV) of the gamble, typically represented as:
@@
where the value $x_{t}$ is realized at time $t$ and is discounted by a factor of $d$, reflecting the time preference of the decision-maker. Note that a risk-neutral and time-consistent agent should always select the option that maximizes EV and PV. However, extensive research in economics and psychology demonstrates that people systematically deviate from this maximizer model (Kahneman & Tversky, 2013; Gigerenzer & Gaissmaier, 2011; Laibson, 1997; Zhu et al., 2020).
Pretrained, off-the-shelf LLMs, such as the GPT and LLaMA series, have demonstrated behavioral similarities to humans in tasks involving risky and intertemporal choices (Horton, 2023; Binz & Schulz, 2023b; Manning et al., 2024). However, the embeddings generated by these LLMs are not by default able to account for human data. For example, embeddings from the LLaMA-1-65B model, when not finetuned on human risky choices, poorly predict those choices (Binz & Schulz, 2023a). Therefore, understanding what enables LLMs to exhibit human-like decision-making behavior remains an unresolved challenge.
We propose a hypothesis about how human-like decision patterns might be produced in language models for risky and intertemporal choice: such human-like behaviors might arise from a language model trained on ecologically valid calculations of expectations. This suggests that human-like biases could result from the training task and numerical reasoning, independent of natural language supervision. As a corollary, this hypothesis also implies that deviations from rational choice in humans could be primarily explained by computational errors during the EV or PV calculations. To test this hypothesis, we generate a series of synthetic datasets that contain expected value computations; examples include 0.5*100=+50, 0.8*1+0.8^2*10=+7.2, and 30*0.79-261*0.83=-192.93. Subsequently, we train a small, randomly-initialized language model (approximately 10M parameters) on these datasets. After training, we extract embeddings from the now pretrained language model and analyze how well they can account for human choices.
@@ 2 Background
The question of whether language models can serve as models of human cognition has sparked intense debate within the fields of machine learning and cognitive science (Frank, 2023b; Messeri & Crockett, 2024; Griffiths et al., 2023; Horton, 2023). Although there are behavioral similarities between off-the-shelf LLMs and humans, these do not inherently qualify LLMs as effective cognitive models (c.f. the Clever Hans effect) (Shiffrin & Mitchell, 2023). There are compelling reasons why current LLMs may not be suitable as cognitive models. First, LLMs are trained on datasets vastly larger than those available to human learners (Frank, 2023a). Second, LLMs may have already been trained on the test questions, particularly if the training data is undisclosed and poorly controlled (Jiang et al., 2024). Third, the inclusion of value alignment steps, such as Reinforcement Learning from Human Feedback (Ziegler et al., 2019) and Direct Preference Optimization (Rafailov et al., 2024), may artificially enhance human-like behaviors in leading LLMs. Finally, the immense size of deep neural networks and the proprietary nature of leading LLMs hinder detailed investigation into their internal representations.
@@
Researchers have taken steps to address some of the concerns associated with using LLMs as cognitive models. One branch of research looks at the potential benefits of fine-tuning off-the-shelf LLMs to better understand human cognition. For instance, fine-tuning the LLaMA-1-65B model (Touvron et al., 2023) on datasets of human choices results in a model that can predict human data more accurately than traditional cognitive models (Binz & Schulz, 2023a). Although the specific mechanisms underlying the finetuned LLM’s ability to replicate human choices remain unclear, this work serves as a proof-of-concept and suggests that the embeddings learned from extensive pretraining on Internet text and/or from the value alignment process may offer valuable insights into human cognitive processes that complement those provided by traditional cognitive models.
@@
Another line of research emphasizes the importance of the composition of synthetic datasets, which are critical for enhancing certain capabilities of LLMs (Lee et al., 2023). These studies typically assess LLMs based on their problem-solving abilities rather than their human-likeness. For example, it has been found that larger language models tend to perform arithmetic tasks, such as addition and multiplication, better than their smaller counterparts (Yuan et al., 2023). Moreover, the order of input and the inclusion of intermediate information about task decomposition have been shown to facilitate chain-of-thought reasoning, which significantly helps smaller language models in mastering arithmetic tasks (Lee et al., 2023).
@@
Finally, there is a precedent for the idea that pretraining machine learning models on synthetic datasets can improve performance in predicting human decisions. Bourgin et al. (2019) showed that a model pretrained on choice data generated from a psychological theory could perform extremely well in predicting human decisions when fine-tuned with a small amount of human data. Our approach builds on this idea, but reduces it to the most primitive components – rather than pretraining on data generated from a psychological theory, we pretrain on a task that captures the basic computations required to make rational decisions.
@@ 3 Arithmetic-GPT: A Small Language Model Trained to Perform Arithmetic
In this paper, we confront the challenges of making language models as cognitive models head-on. First, we define a data generation algorithm to produce synthetic datasets, thereby gaining complete control over the training data for language models and addressing issues related to data gaps and contamination. Second, we have direct access to the neural activation patterns that are crucial for decision-making processes. This approach allows us to more thoroughly evaluate and understand the capabilities and limitations of language models.
@@ 3.1 Model Details
Our small language model employs a standard architecture for a Generative Pretrained Transformer (GPT) model (Vaswani et al., 2017; Radford et al., 2019). Detailed specifications of the model architecture are provided in Table 1.
@@ 3.2 Synthetic Datasets
At the heart of the EV and PV computations lies the multiplication of two real numbers. Typically, each outcome appears in the computation as either a probability multiplied by a value, or a discount factor multiplied by a value. Probabilities are real numbers ranging from 0 to 1, represented with a precision of up to two decimal places. Similarly, values are real numbers that range from 0 to 300, also with a maximum of two decimal places. We selected these ranges to align with the numerical scope of human experimental studies. A single training example involves either the addition or subtraction of two simulated outcomes, which together constitute the left-hand side of an equation. We then compute the corresponding result and place it on the right-hand side of the equation. In total, we randomly simulated 1M such equations.
In our experiment, we evaluate four variants of the data generation algorithm by manipulating the frequency of probabilities and values (see Table 3). The Uniform synthetic data generates probabilities and values with maximum uncertainty; probabilities are uniformly distributed between 0 and 1 (i.e., $U[0,1]$), and values range from 0 to 300 (i.e., $U[0,300]$). Conversely, the Ecological synthetic data generates probabilities following a Beta distribution $\text{Beta}(0.27,0.27)$ (Zhu et al., 2020) and values according to a power-law distribution with an exponent of $-0.945$ for the same range (Stewart et al., 2006). These distributions are chosen because they have been shown to closely match the natural frequencies of probabilities and values in real-world scenarios (Stewart et al., 2006; Zhu et al., 2020) (see Figure 1B for details). For both uniform and ecological synthetic datasets, we also created a matching dataset where the answers on the right-hand side are generated with a 50% chance of displaying the incorrect sign (i.e., the ablated variants in Table 3). In each of the four synthetic datasets, we randomly masked 10% of the probability values using a special <AMB> token to denote unknown probabilities.
@@ 3.3 Pretraining Details
We trained our Arithmetic-GPT models from scratch with a context length of 26. Batch size was set to 2048 lines of mathematical equations, randomly sampled from the synthetic datasets, and the learning rate was $10^{-3}$. To optionally terminate the training process, we designated 90% of the synthetic dataset as the training set, reserving the remaining 10% for validation. We used cross-entropy loss to measure the discrepancy between the predicted and target sequences. Training was stopped when the validation loss plateaued, with validation loss evaluated every 100 epochs (see Figure A1 in Appendix C for an example learning curve). The AdamW optimizer was used.
@@ 3.4 Human Targets
To investigate whether Arithmetic-GPT contains information pertinent to explaining human decision-making, we reanalyzed recent experiments in which people were asked to make risky and intertemporal choices (Peterson et al., 2021; Erev et al., 2017; Gershman & Bhui, 2020; Agrawal et al., 2023). The primary reason for examining these particular types of human choices is that calculations of expected value are integral to making rational decisions (see Equations 1 and 2). In other words, they are computationally equivalent tasks under the assumption of rationality.
As summarized in Table 2, we sampled four existing datasets from the literature, which included two large-scale experiments (Peterson et al., 2021; Agrawal et al., 2023). In experiments involving risky choices (Peterson et al., 2021; Erev et al., 2017), participants were often presented with two options, each fully describing the details of a gamble (see Figure 1A risky choices). In cases involving ambiguous gambles where probabilities are unknown, we used the special token <AMB> to denote gambles with unknown probabilities. We excluded decision-with-feedback trials, as these would require additional assumptions about how individuals respond to feedback (but see Table A2 of Appendix B for a comparison based on the entire choices13k dataset). In intertemporal choice tasks (Gershman & Bhui, 2020; Chávez et al., 2017; Agrawal et al., 2023), participants were also presented with two options, typically offering a choice between a smaller, sooner payoff and a larger, later payoff (see Figure 1A intertemporal choices). Without loss of generality, we fixed the annual discount factor at $d_{\text{year}}=0.85$ throughout the paper. This also corresponds to a monthly discount factor of $d_{\text{month}}=0.98$ and a daily discount factor of $d_{\text{day}}=0.99$. We rescaled the values in both options by the same factor to fit within the specified range.
@@
Peterson et al. (2021) choices13k Risky choices 15,153 13,006
Erev et al. (2017) cpc18 Risky choices 446 270
Gershman & Bhui (2020) gershman20 Intertemporal choices 221 4,794
Agrawal et al. (2023) agrawal23 Intertemporal choices 12,906 9,853
Note. In our analysis of the choices13k and cpc18 datasets, we excluded trials involving risky choices made with feedback. Modeling how individuals respond to feedback requires additional cognitive mechanisms beyond the scope of this work.
@@ 4 Other Models
We also conduct a model comparison on human data, evaluating the following approaches: (i) classical behavioral models such as prospect theory (Kahneman & Tversky, 2013) and the hyperbolic discounting model (Laibson, 1997); (ii) a neural network directly trained on the human datasets; (iii) an untrained Arithmetic-GPT model; and (iv) an off-the-shelf, open-weight LLM, LLaMA-3-70B-Instruct. [^2] https://llama.meta.com/llama3/
Classical behavioral models. To explain human risky choices, prospect theory proposes an S-shaped utility function that is concave for gains, convex for losses, and steeper for losses than for gains (reflecting loss aversion). The utility function is represented mathematically as:
@@
Contrary to the consistent risk preferences implied by Equation 1 or other monotonic transformations of value, prospect theory suggests that risk preferences are inconsistent across different values and probabilities. This inconsistency results in incoherent choices when individuals are faced with risky decisions (Tversky & Kahneman, 1992).
@@
where $x_{t}$ is the value to be received at future time $t$, and $k$ is the discount factor that quantifies the degree of time preference. In contrast to the consistent time preferences implied by Equation 2, the hyperbolic discounting model suggests that time preferences are inconsistent across different time horizons, leading to stronger preference to immediate over future rewards (Laibson, 1997).
MLPs directly trained on human choices. We also implemented Multilayer Perceptrons (MLPs) with a single hidden layer containing 320 neurons and using the sigmoid activation function. These MLPs were directly trained on each of the four human datasets, using the stimuli features as input to predict choice rates. These MLPs potentially capture an upper bound of the explainable variance within human data (Agrawal et al., 2020). However, they may have overlooked significant constraints from the original psychological experiments, as these models use task features as input rather than the actual stimuli presented in texts or the stylized representations required for rational agents, as in our Arithmetic-GPT.
@@
Given these features of LLaMA3, we presented each choice problem in two different formats: text-based and arithmetic equation-based. The text format converts each choice problem into a descriptive narrative, simulating the stimuli presented to human participants (see Appendix A for a detailed description of the text prompts). We instructed the model to report its selection between the two options, using the log probability of the chosen option to determine the model’s predicted choice rates for the corresponding option. In contrast, the arithmetic-equation format presents the choice problems as a series of arithmetic computations required by a rational agent. Note that this format is identical to the input used for Arithmetic-GPT. We obtained the embeddings from LLaMA3 for each choice problem represented in the arithmetic-equation format.